# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v1

@@ 2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
$M_{\mathrm{cross}}(T):=B\,C_{IW}(T)\,\big(\Sigma^{1/2}\big)^{\top}+\Sigma^{1/2}\,C_{IW}(T)^{\top}\,B^{\top}\;\in\;\mathbb{R}^{d\times d}.$ (50)
Then the Gaussian-$q$ decision-time reference under constant portfolios is characterized by
$\Big(\gamma T\Sigma+(\gamma-1)\big(B\,C_{I}(T)\,B^{\top}+M_{\mathrm{cross}}(T)\big)\Big)\,\pi_{q,\gamma}^{\mathrm{const}}(T)=B\,m_{I}(T),\hskip 18.49988pt(\gamma>1),$ (51)
equivalently
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\Big(\gamma T\Sigma+(\gamma-1)\big(B\,C_{I}(T)\,B^{\top}+M_{\mathrm{cross}}(T)\big)\Big)^{-1}B\,m_{I}(T).$
When $\rho=0$, we have $C_{IW}(T)=0$ and $M_{\mathrm{cross}}(T)=0$, recovering the independence-case shrinkage reference.
2.4 Why dynamic programming and deep PDE surrogates break down in high-dimensional uncertain markets
This subsection explains why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as practical baselines in the high-dimensional uncertain markets targeted here. DP is conceptually sound in low-dimensional Markovian settings (fleming2006controlled; pham2009continuous), but two issues dominate in our regime: (i) numerically learning the value-function derivatives required for optimal policies becomes prohibitive as dimension and nonlinearity grow, and (ii) principled parameter uncertainty magnifies these difficulties.
Classical HJB: curse of dimensionality and full nonlinearity. With deterministic parameters, DP leads to an HJB for $V(t,x,y)$ (fleming2006controlled). Grid-based solvers scale exponentially in the state dimension (bellman1961adaptive; kushner2001numerical). In portfolio problems with $d$ assets and $m$ factors, the natural state already has dimension $m+2$, so even modest discretizations require $N^{m+2}$ grid points. Moreover, constraints, transaction costs, and non-affine dynamics typically yield fully nonlinear HJBs, where stable monotone schemes are delicate even in moderate dimension and become impractical in the regime we target (kushner2001numerical).
Deep PDE surrogates: fewer grids, same derivative bottleneck. PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (raissi2019physics; sirignano2018dgm; han2018solving; beck2019machine), but for fully nonlinear portfolio HJBs they remain value-function-based: they must implicitly learn high-dimensional gradients/Hessians and, crucially, mixed sensitivities (e.g. $V_{xy}$) that drive intertemporal hedging. In practice this induces nonconvex, ill-conditioned objectives (due to control suprema and nonlinear derivative dependence) and training signals that do not reliably control the specific derivative components needed for stable hedging demands in high dimension.
Latent parameter uncertainty: belief-state blowup and $\theta$-blind aggregation. A principled DP treatment augments the state with a posterior/belief over parameters, leading to a value function $V(t,x,y,\Pi)$ on a space of measures in general (bensoussan1985optimal; pham2017dynamic). Even when finite-dimensional conjugate summaries exist, the enlarged HJB is substantially harder. For deep surrogates, uncertainty either requires solving many $\theta$-conditional problems (expensive) or treating $\theta$ as an extra input (higher effective dimension, worse conditioning). In our deployable $\theta$-blind setting, a single policy must perform well under $\theta\sim q$, coupling heterogeneous models and potentially causing high-variance gradients and cancellation across parameter draws.
We therefore avoid value-function PDE/BSDE baselines in this regime and instead work with a $q$-aggregated Pontryagin stationarity condition and projection map, estimating expectations over $\theta\sim q$ via Monte Carlo inside the simulator. While $\theta$-conditional PMP objects can still be computed under frozen-$\theta$ simulations for inspection, our deployable target and guarantees are expressed in terms of $q$-aggregated stationarity, motivating the simulation-based methods in Section 3.
@@ 3 Pontryagin–Guided Policy Optimization under Latent Parameter Uncertainty
We study the fixed-$q$ ex–ante portfolio choice problem of Section 2 under latent parameter uncertainty $\theta\sim q$. The investor must deploy a $\theta$-blind policy (Remark 1), so the control can depend on observable states $(t,X_{t},Y_{t})$ but cannot take $\theta$ as an input. We restrict attention to Markov feedback policies parameterized by a neural network $\pi_{\varphi}(t,x,y)$.
Our solution approach follows a two-stage pipeline:
  • Stage 1 (PG–DPO). We perform stochastic gradient ascent on the ex–ante objective $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$, sampling $\theta$ only inside the simulator while keeping $\pi_{\varphi}$ deployable and $\theta$-blind.
  • Stage 2 (Pontryagin projection). Under a warm-up policy, we estimate Pontryagin sensitivity objects by BPTT (conditionally on frozen $\theta$), aggregate them across $\theta\sim q$, and construct a single deployable portfolio by projecting onto the aggregated first-order condition (20).
A practical subtlety is that the $q$-aggregated Pontryagin condition involves mixed moments across $\theta$ (products of $\theta$-dependent costates and $\theta$-dependent coefficients). In moderate to high dimensions, these quantities can be statistically noisy under finite Monte Carlo budgets. In our implementation, the main stabilization mechanisms are (i) estimating stage 2 objects under a warm-up policy (two-time-scale stabilization), (ii) computing the same projection in a residual/control-variate form (Section 3.3.1), and (iii) amortizing projection via interactive distillation (Section 3.3.2).
Section 3.1 reviews baseline PG–DPO and the conditional BPTT–PMP correspondence. Section 3.2 develops the stage 2 $q$-aggregated projection under latent $\theta$, together with a residual-based policy-gap guarantee. Section 3.3 records two practical couplings between stage 1 and stage 2 (residual form and interactive distillation).
3.1 PG–DPO as stochastic gradient ascent and BPTT–PMP correspondence
@@
which does not take $\theta$ as an input. For a fixed frozen $\theta$, the $\theta$-conditional state is $(X_{t}^{\pi,\theta},Y_{t}^{\theta})_{t\in[0,T]}$ and evolves as
@@
$dY_{t}^{\theta}$ $=a\big(Y_{t}^{\theta},\theta\big)\,dt+\beta\big(Y_{t}^{\theta},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m}.$ (54)
For each fixed $\theta$ we evaluate $\pi_{\varphi}$ by the conditional objective
@@
where the last expectation is joint over $\theta\sim q$ and $(W,W^{Y})$. Thus $\sup_{\varphi}J(\varphi)$ is a standard stochastic optimization problem: $\theta$ is sampled inside the simulator while the policy remains $\theta$-blind.
@@
We discretize $[0,T]$ into $N$ steps of length $\Delta t$ and approximate (53)–(54) by an Euler scheme. For episode $i$ we denote the discrete state by $(X_{k}^{(i)},Y_{k}^{(i)})_{k=0,\dots,N}$ and write $\theta^{(i)}$ for the frozen parameter used to generate that simulated environment. Given $\pi_{\varphi}$ and Brownian increments, the mapping
@@
is a finite computational graph, so automatic differentiation computes exact discrete gradients $\nabla_{\varphi}U(X_{N}^{(i)})$.
A typical PG–DPO update samples a mini-batch of initial states $\{(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})\}_{i=1}^{M}$ from a user-chosen training distribution $\nu$ on $0,T)\times(0,\infty)\times\mathbb{R}^{m}$, samples $\theta\sim q$ inside the simulator (unseen by the policy) and holds it frozen for the update, and simulates forward:
$Y_{k+1}^{(i)}$ $=Y_{k}^{(i)}+a\big(Y_{k}^{(i)},\theta\big)\Delta t^{(i)}+\beta\big(Y_{k}^{(i)},\theta\big)\Delta W_{k}^{Y,(i)},$
$X_{k+1}^{(i)}$ $=X_{k}^{(i)}+X_{k}^{(i)}\Big(r+\pi_{\varphi}\big(t_{k}^{(i)},X_{k}^{(i)},Y_{k}^{(i)}\big)^{\top}b\big(Y_{k}^{(i)},\theta\big)\Big)\Delta t^{(i)}$
$\hphantom{={}}{}+X_{k}^{(i)}\,\pi_{\varphi}\big(t_{k}^{(i)},X_{k}^{(i)},Y_{k}^{(i)}\big)^{\top}\sigma\big(Y_{k}^{(i)},\theta\big)\Delta W_{k}^{(i)},$
starting from $X_{0}^{(i)}=x_{0}^{(i)}$, $Y_{0}^{(i)}=y_{0}^{(i)}$. The episode reward is
@@
Sampling $\theta$ independently per episode (i.e. $\theta^{(i)}\sim q$) or sampling one $\theta\sim q$ per update and reusing it across the batch both yield unbiased stochastic gradients for $J(\varphi)$.
Pathwise costates from BPTT and the (conditional) BPTT–PMP correspondence.
BPTT returns not only $\nabla_{\varphi}J^{(i)}(\varphi)$ but also pathwise sensitivities with respect to intermediate state variables, which coincide with discrete-time adjoint variables (costates) in the sense of Pontryagin. For a single episode (suppressing $i$ and $\theta$ in notation), define the pathwise wealth costate
@@
and the additional pathwise sensitivity objects used in projected controls:
$p_{x,k}:=\frac{\partial p_{k}}{\partial X_{k}},\hskip 18.49988ptp_{y,k}:=\frac{\partial p_{k}}{\partial Y_{k}},\hskip 18.49988ptk=0,\dots,N.$ (60)
Theorem 2 (BPTT–PMP correspondence (conditional on $\theta$, uniform on compacts)).
Fix $\theta\in\Theta$ and assume standard regularity conditions ensuring (i) well-posedness of the $\theta$-conditional forward SDE (53)–(54) under the $\theta$-blind policy $\pi_{\varphi}$ and (ii) well-posedness of the associated $\theta$-conditional stochastic maximum principle (adjoint) system. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the resulting continuous-time Pontryagin objects under $\pi_{\varphi}$ (and, in smooth Markov regimes, the corresponding spatial derivatives of the decoupling field). Let $(p_{k},p_{x,k},p_{y,k})$ be the discrete pathwise quantities computed by BPTT for the Euler discretization with step $\Delta t$, as defined in (59)–(60).
Then, as $\Delta t\to 0$, the BPTT-induced discrete adjoints converge to their continuous-time counterparts in an appropriate mean-square sense (along trajectories). Moreover, for any compact set $K\subset\Theta$, the constants in the convergence bounds can be chosen uniformly for all $\theta\in K$.
@@
3.2 Projected PG–DPO under latent $\theta$: $q$-aggregated projection and a residual-based policy-gap bound
Stage 2 is a projection step: given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ (from stage 1), we estimate $\theta$-conditional Pontryagin sensitivity objects by BPTT/Monte Carlo under frozen $\theta\sim q$, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated Pontryagin stationarity condition derived in Section 2.2. The main point is that the aggregated first-order condition is affine in the portfolio control; hence it induces a statewise linear system and, on a suitable working domain, a concrete projection map from estimated Pontryagin objects to a portfolio rule.
@@
Fix a query state $z=(t,x,y)\in D$ and a frozen parameter $\theta$. We simulate trajectories under $\pi^{\mathrm{warm}}$ and compute pathwise Pontryagin sensitivity objects by autodiff/BPTT; averaging over $M_{\mathrm{MC}}$ trajectories yields Monte Carlo estimates
$\widehat{p}_{t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{x,t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{y,t}^{\theta}(z).$ (62)
@@
$\widehat{G}_{t}^{\theta}(t,x,y)$ $:=\widehat{p}_{t}^{\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,\widehat{p}_{y,t}^{\theta}(t,x,y)\in\mathbb{R}^{d}.$ (64)
Aggregating across $\theta\sim q$ (approximated in practice by sampling $M_{\theta}$ frozen parameters) gives
@@
To connect the projected policy (67) to a locally optimal deployable $\theta$-blind policy, we measure how well the warm-up policy satisfies the population mixed-moment aggregated stationarity. Let $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ denote the mixed-moment $q$-aggregated projection inputs induced by a policy $\pi$ (i.e. the objects in (23) evaluated using the $\theta$-conditional Pontryagin objects generated by $\pi$). Define the warm-up aggregated stationarity residual on $D$ by
@@
In practice we monitor the estimator $\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{G}_{t}^{\mathrm{mix}}$ computed from the same BPTT/Monte Carlo pipeline.
A technical point is that a global small-gain condition of the form $C_{1}<1$ can be overly restrictive. Following the slab-wise philosophy in our prior PGDPO analysis (e.g. huh2025breaking), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab. Concretely, assume $D$ carries a time coordinate and fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$. Let

### v2

@@ 2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
$M_{\mathrm{cross}}(\tau):=B\,C_{IW}(\tau)\,\big(\Sigma^{1/2}\big)^{\top}+\Sigma^{1/2}\,C_{IW}(\tau)^{\top}\,B^{\top}\;\in\;\mathbb{R}^{d\times d}.$ (42)
Then the Gaussian decision-time reference under constant portfolios is characterized by
$\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=B\,m_{I}(t,\tau),\hskip 18.49988pt(\gamma>1).$ (43)
Equivalently,
$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)^{-1}B\,m_{I}(t,\tau).$
When $\rho=0$, we have $C_{IW}(\tau)=0$ and $M_{\mathrm{cross}}(\tau)=0$, and the reference reduces to the independence-case shrinkage form obtained by plugging the effective Gaussian law (40) into the static Gaussian formula of Section 2.3.1:
$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{\bar{\theta}}(t,\tau)\big)^{-1}m_{\bar{\theta}}(t,\tau),\hskip 18.49988pt(\rho=0,\ \gamma>1).$
@@ 3 Pontryagin–Guided Policy Optimization under Latent Parameter Uncertainty
We solve the fixed-$q$ ex–ante portfolio problem of Section 2 under a latent, time-constant parameter $\theta\sim q$. The deployed controller is $\theta$-blind: it depends only on observable states $(t,X_{t},Y_{t})$ and never conditions on the realized $\theta$. Uncertainty enters only through sampling $\theta$ inside the simulator.
Our solver has two components. Stage 1 (PG–DPO) performs stochastic gradient ascent on $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$ by differentiating through an Euler simulator via BPTT. Stage 2 (P–PGDPO) aggregates the BPTT costate blocks across $\theta\sim q$ and enforces the $q$-aggregated Pontryagin first-order condition, yielding a deployable projected control.
Section 3.1 establishes the conditional BPTT–PMP correspondence and specifies the adjoint blocks needed for projection. Section 3.2 develops the $q$-aggregated projection and a residual-based policy-gap bound. Practical coupling mechanisms (residual-form projection and interactive distillation) are deferred to Appendix F.
3.1 PG–DPO as stochastic gradient ascent and a conditional BPTT–PMP correspondence
@@
which does not take $\theta$ as an input. Conditional on a fixed frozen $\theta$, the controlled state $(X_{t}^{\pi,\theta},Y_{t}^{\theta})_{t\in[0,T]}$ evolves as
@@
$dY_{t}^{\theta}$ $=a\big(Y_{t}^{\theta},\theta\big)\,dt+\beta\big(Y_{t}^{\theta},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m},$ (46)
where $(W,W^{Y})$ may be instantaneously correlated as in Section 2. For each fixed $\theta$ we evaluate $\pi_{\varphi}$ by the conditional objective
@@
where the last expectation is joint over $\theta\sim q$ and $(W,W^{Y})$. Thus $\sup_{\varphi}J(\varphi)$ is a stochastic optimization problem in which $\theta$ is sampled inside the simulator while the policy remains $\theta$-blind.
@@
We discretize the horizon using an Euler scheme. In training we may sample initial states $z_{0}^{(i)}=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})$ from a user-chosen distribution $\nu$ on $0,T)\times(0,\infty)\times\mathbb{R}^{m}$. Given $t_{0}^{(i)}$, we discretize $t_{0}^{(i)},T]$ into $N$ steps with step size
$\Delta t^{(i)}:=\frac{T-t_{0}^{(i)}}{N},\hskip 18.49988ptt_{k}^{(i)}:=t_{0}^{(i)}+k\Delta t^{(i)},\qquad k=0,\dots,N.$
For episode $i$, we denote the discrete state by $(X_{k}^{(i)},Y_{k}^{(i)})_{k=0,\dots,N}$ and write $\theta^{(i)}$ for the frozen parameter used to generate that simulated environment. Given $\pi_{\varphi}$ and Brownian increments, the mapping
@@
is a finite computational graph; automatic differentiation thus computes exact discrete-time gradients $\nabla_{\varphi}U(X_{N}^{(i)})$ for the chosen discretization.
A typical PG–DPO update samples a mini-batch $\{z_{0}^{(i)}\}_{i=1}^{M}\sim\nu$, samples latent parameters either independently per episode ($\theta^{(i)}\sim q$) or shared within the batch (one $\theta\sim q$ reused across $i$), simulates Euler rollouts, and performs stochastic gradient ascent. Suppressing episode indices and writing $\theta$ for the frozen draw used in the rollout, the Euler recursion is
$Y_{k+1}$ $=Y_{k}+a\big(Y_{k},\theta\big)\Delta t+\beta\big(Y_{k},\theta\big)\Delta W_{k}^{Y},$
$X_{k+1}$ $=X_{k}+X_{k}\Big(r+\pi_{\varphi}\big(t_{k},X_{k},Y_{k}\big)^{\top}b\big(Y_{k},\theta\big)\Big)\Delta t+X_{k}\,\pi_{\varphi}\big(t_{k},X_{k},Y_{k}\big)^{\top}\sigma\big(Y_{k},\theta\big)\Delta W_{k},$
starting from $(t_{0},X_{0},Y_{0})=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})$. The episode reward is
@@
Under standard interchange conditions (dominated convergence / uniform integrability for the discretized simulator), this is an unbiased stochastic gradient for the discretized ex–ante objective; as $\Delta t\to 0$ it approximates the continuous-time objective.
BPTT adjoints and their interpretation: first adjoint and second-adjoint blocks.
BPTT returns not only $\nabla_{\varphi}J^{(i)}(\varphi)$ but also adjoint variables with respect to intermediate state variables. For a single episode (suppressing $i$ and $\theta$ in notation), define the pathwise wealth adjoint
@@
which is the discrete-time analogue of the wealth-component first adjoint. For the Pontryagin projection in Section 3.2, we also require curvature-type objects. In smooth Markov regimes these correspond to $(V_{xx},V_{xy})$; in the stochastic maximum principle they are naturally interpreted as blocks of the second adjoint. Accordingly, we track the following discrete-time blocks (computed by higher-order automatic differentiation):
$p_{x,k}:=\frac{\partial p_{k}}{\partial X_{k}}\;=\;\frac{\partial^{2}U(X_{N})}{\partial X_{k}^{2}},\hskip 18.49988ptp_{y,k}:=\frac{\partial p_{k}}{\partial Y_{k}}\;=\;\frac{\partial^{2}U(X_{N})}{\partial Y_{k}\,\partial X_{k}},\hskip 18.49988ptk=0,\dots,N.$ (52)
(Here $p_{y,k}\in\mathbb{R}^{m}$ is the wealth–factor cross block.) These are the discrete-time objects that enter the affine-in-control Hamiltonian gradient and hence the projection step.
Pathwise vs. SMP-adapted adjoints. The quantities in (51)–(52) are pathwise (sample-wise) derivatives on the simulated computation graph and, for $k<N$, they need not be adapted to time-$t_{k}$ information (they can depend on future increments). To connect them to the stochastic maximum principle (which yields adapted adjoint processes and associated martingale terms), one performs the standard conditional $L^{2}$ projection onto the span of the driving increments on each step, yielding a discrete-time BSDE representation. Concretely, for a fixed frozen $\theta$ and the simulator filtration
$G_{t_{k}}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t_{k}\}\big)\qquad\text{(with the usual augmentation)},$
we define the adapted (projected) discrete adjoints
$p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid G_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid G_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid G_{t_{k}}^{\theta}].$
Theorem 2 concerns the convergence of these adapted discrete objects (together with the associated one-step BSDE martingale coefficients) to the continuous-time Pontryagin adjoints.
Theorem 2 (BPTT–PMP correspondence (conditional on $\theta$, including second-adjoint blocks; uniform on compacts)).
Fix $\theta\in\Theta$ and assume standard regularity conditions ensuring (i) well-posedness of the $\theta$-conditional forward SDE (45)–(46) under the $\theta$-blind policy $\pi_{\varphi}$, and (ii) well-posedness of the associated $\theta$-conditional stochastic maximum principle, including the wealth-component first adjoint and the relevant blocks of the second adjoint needed for the portfolio Hamiltonian gradient. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote these continuous-time Pontryagin objects under $\pi_{\varphi}$. In smooth Markov regimes, they coincide (along the controlled state process) with $(V_{x},V_{xx},V_{xy})$ for the $\theta$-conditional value function.
Let $(p_{k},p_{x,k},p_{y,k})$ be the discrete pathwise quantities computed by BPTT for the Euler discretization with step $\Delta t$, as defined in (51)–(52). Let $(p_{k}^{\Delta t,\theta},p_{x,k}^{\Delta t,\theta},p_{y,k}^{\Delta t,\theta})$ be their adapted projections onto $G_{t_{k}}^{\theta}$ as defined above, and let the associated discrete-time BSDE martingale coefficients be defined via the one-step conditional $L^{2}$ projection on each step. Then, as $\Delta t\to 0$, the projected (adapted) BPTT-induced discrete adjoints converge to their continuous-time counterparts in an appropriate mean-square sense (along trajectories). Moreover, for any compact set $K\subset\Theta$, the constants in the convergence bounds can be chosen uniformly for all $\theta\in K$.
@@
3.2 Projected PG–DPO under latent $\theta$: $q$-aggregated projection, diagnostics, and a residual-based policy-gap bound
Stage 2 is a projection step. Given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ (from stage 1), we estimate $\theta$-conditional Pontryagin adjoint objects under frozen $\theta\sim q$ by BPTT/Monte Carlo, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated stationarity condition derived in Section 2.2. Equivalently, Stage 2 can be viewed as a plug-in approximation of one application of the population projection map $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ (defined below) evaluated at $\pi=\pi^{\mathrm{warm}}$ on the working domain. The key point is that the aggregated first-order condition is affine in the portfolio control; hence it induces a statewise linear system and, on a suitable working domain, a concrete projection map from (estimated) adjoint objects to a portfolio rule.
Throughout, the adjoint objects used in projection consist of the wealth-component first adjoint together with the relevant blocks of the second adjoint (cf. Section 3.1): in smooth Markov regimes these correspond to $(V_{x},V_{xx},V_{xy})$ evaluated along the controlled state.
@@
Fix a query state $z=(t,x,y)\in D$ and a frozen parameter $\theta$. We simulate trajectories under $\pi^{\mathrm{warm}}$ and compute discrete-time adjoint blocks by autodiff/BPTT on the Euler simulator; averaging over $M_{\mathrm{MC}}$ trajectories yields Monte Carlo estimates
$\widehat{p}_{t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{x,t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{y,t}^{\theta}(z),$ (54)
where $\widehat{p}_{t}^{\theta}$ corresponds to the wealth-component first adjoint and $(\widehat{p}_{x,t}^{\theta},\widehat{p}_{y,t}^{\theta})$ correspond to the blocks of the second adjoint relevant for the Hamiltonian gradient. In the analysis, these are interpreted as estimators of the corresponding $\theta$-conditional Pontryagin objects (cf. the conditional BPTT–PMP correspondence in Section 3.1).
@@
$\widehat{G}_{t}^{\theta}(t,x,y)$ $:=\widehat{p}_{t}^{\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,\widehat{p}_{y,t}^{\theta}(t,x,y)\in\mathbb{R}^{d},$ (56)
where $\Sigma(\cdot,\theta)$ and $\Sigma_{SY}(\cdot,\theta)$ are the instantaneous covariance objects appearing in the $\theta$-conditional Hamiltonian (as in Section 2).
Aggregating across $\theta\sim q$ (approximated in practice by sampling $M_{\theta}$ frozen parameters) gives the mixed-moment estimators
@@
Stabilized inversion and projection diagnostics (implementation). In practice, finite-sample estimates can make $\widehat{A}_{t}$ ill-conditioned or even singular on parts of $D$. Our default implementation therefore uses standard stabilizers (Appendix E.5), including: (i) ridge regularization $\widehat{A}_{t,\lambda}:=\widehat{A}_{t}+\lambda I_{d}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\lambda}:=-\widehat{A}_{t,\lambda}^{-1}\widehat{G}_{t}^{\mathrm{mix}}$, (ii) condition-number or eigenvalue diagnostics to skip/clip projection on unreliable states (falling back to $\pi^{\mathrm{warm}}$), and (iii) residual-form projection (Section F.1) to reduce variance. The analysis below is stated for the population projection $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ and for small estimator perturbations, which is precisely the regime enforced by these stabilizers.
Remark (decoupled aggregation heuristic). The mixed-moment aggregation (57)–(58) is the object appearing in the exact $q$-aggregated first-order condition. In some high-dimensional experiments, we additionally consider a variance-reduction heuristic that replaces mixed moments by products of empirical means, e.g. $\mathbb{E}_{\theta}[\widehat{p}_{t}^{\theta}\,b(y,\theta)]\approx\mathbb{E}_{\theta}[\widehat{p}_{t}^{\theta}]\cdot\mathbb{E}_{\theta}[b(y,\theta)]$, and similarly for the other terms. All theoretical statements in this subsection are for the mixed-moment projection (59).
@@
To connect (59) to a locally optimal deployable $\theta$-blind policy, we measure how well the warm-up policy satisfies the population mixed-moment aggregated stationarity. Let $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ denote the mixed-moment $q$-aggregated projection inputs induced by a policy $\pi$ (i.e. the objects in (18) evaluated using the $\theta$-conditional Pontryagin adjoint blocks generated by $\pi$). Define the warm-up aggregated stationarity residual on $D$ by
@@
Note that $r^{\mathrm{warm}}_{\mathrm{FOC,mix}}=0$ is equivalent to $\pi^{\mathrm{warm}}$ being a fixed point of $T$ on $D$. More generally, the population projection admits the residual form
$T(\pi^{\mathrm{warm}})\;=\;\pi^{\mathrm{warm}}\;-\;A_{\pi^{\mathrm{warm}}}^{-1}\,r^{\mathrm{warm}}_{\mathrm{FOC,mix}}.$ (61)
In practice we monitor the empirical analogue $\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{G}_{t}^{\mathrm{mix}}$, and use it as a diagnostic to flag states where projection is unreliable (cf. Appendix E.5). (Equivalently, $\widehat{\pi}^{\mathrm{agg,mix}}=\pi^{\mathrm{warm}}-\widehat{A}_{t}^{-1}\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}$, or its stabilized variant, which motivates the residual-form implementation.)
BPTT/Monte Carlo error term (one explicit choice). To make the discretization/Monte Carlo dependence in (65) explicit, one may take for example
$\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}):=\big\|\widehat{A}_{t}-A_{\pi^{\mathrm{warm}}}\big\|_{L^{2}(\mu)}+\big\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}\big\|_{L^{2}(\mu)}.$ (62)
Under standard Lipschitz/bounded-moment assumptions (Euler strong error and Monte Carlo sampling error), one typically has a scaling of the form $\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta})\lesssim\Delta t^{1/2}+M_{\mathrm{MC}}^{-1/2}+M_{\theta}^{-1/2}$, up to constants depending on the working domain and stability bounds.
A technical point is that a global small-gain condition of the form $C_{1}<1$ can be overly restrictive. Following the slab-wise philosophy in our prior PGDPO analysis (e.g. Huh et al. (2025b, Appendix B)), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab. Concretely, assume $D$ carries a time coordinate and fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$. Let